# TLDR: Unsupervised Goal-Conditioned RL via Temporal Distance-Aware Representations

### v1

@@ 3.5 Learning Exploration Policy
Similar to LEXA [10], we alternate between goal-reaching episodes and exploration episodes. For goal-reaching episodes, we execute the goal-conditioned policy until the end of the episodes. For exploration episodes, we sample the timestep $T_{G}\sim\text{Unif}(0,T-1)$ at the beginning of each episode and execute the exploration policy if timestep $t\geq T_{G}$.
@@ 3.6 Learning Goal-Conditioned Policy
(g) Quadruped (Pixel)
(h) Kitchen (Pixel)
Figure 3: Benchmark environments. We evaluate our method on $8$ robotic locomotion and manipulation environments.
4 Experiments
In this paper, we propose TLDR, a novel unsupervised GCRL method that utilizes temporal distance-aware representations for both exploration and optimizing a goal-conditioned policy. Through our experiments, we aim to answer the following $3$ questions: (1) Does TLDR explore better compared to other exploration methods? (2) Is our goal-conditioned policy better than prior unsupervised GCRL methods? (3) How crucial is TLDR for goal-conditioned policy learning and exploration?
4.1 Experimental Setup
(a) Ant
(b) HalfCheetah
(c) Humanoid-Run
(d) Quadruped-Escape
(e) AntMaze-Large
(f) AntMaze-Ultra
Figure 4: State coverage on state-based environments. We measure the state coverage of unsupervised exploration methods. Our method consistently shows superior state coverage compared to other methods, except in HalfCheetah compared against METRA.
@@ 4.1 Experimental Setup
We evaluate our method in $6$ state-based environments and $2$ pixel-based environments, as illustrated in Figure 3. For state-based environments, we use Ant and HalfCheetah from OpenAI Gym [31], Humanoid-Run and Quadruped-Escape from DeepMind Control Suite (DMC) [32], AntMaze-Large from D4RL [33], and AntMaze-Ultra [34]. For Humanoid-Run and Quadruped-Escape, we include the 3D coordinates of the agents in their observations. For pixel-based environments, we use Quadruped (Pixel) from METRA [1] and Kitchen (Pixel) from D4RL [33], with the image size of $64\times 64\times 3$ as the observation.
@@
  • METRA [1]: the state-of-the-art unsupervised skill discovery method which leverages temporal distance-aware representations.
  • PEG [2]: the state-of-the-art unsupervised GCRL method which plans to obtain goals with maximum exploration rewards.
@@
Evaluation setups.
Following METRA [1] and PEG [2], we evaluate unsupervised exploration using state coverage or queue state coverage, and evaluate goal-reaching performance using goal distance or the number of reached goals (achieved tasks). State coverage is calculated as the number of $1\times 1$ sized $(x,y)$-bins ($x$-bins for HalfCheetah) occupied by any of the training trajectories. Queue state coverage for Kitchen (Pixel) is the number of tasks achieved at least once during the last $100,000$ environment steps. For Ant, HalfCheetah, Humanoid-Run, and Quadruped (Pixel), we compute the goal distance by randomly selecting a target goal, executing the goal-reaching policy, and measuring the distance between the final state of the policy and the target goal. For AntMaze and Kitchen (Pixel), we measure the number of reached goals and achieved tasks, respectively. More experimental details are described in Appendix A.
@@ 4.2 Quantitative Results
(a) Ant ($\downarrow$)
(b) HalfCheetah ($\downarrow$)
(c) Humanoid-Run ($\downarrow$)
(d) AntMaze-Large ($\uparrow$)
(e) AntMaze-Ultra ($\uparrow$)
Figure 5: Goal-reaching metrics of a goal-conditioned policy. We first report the average distance between goals and the last states of trajectories (lower is better $\downarrow$) in (a) Ant, (b) HalfCheetah, and (c) Humanoid-Run. TLDR achieves a comparable average goal distance to METRA. For AntMaze environments, we report the number of pre-defined goals reached by a goal-reaching policy ($7$ for (d) AntMaze-Large and $21$ for (e) AntMaze-Ultra), and TLDR significantly outperforms prior works.
In Figure 4, we compare the state coverage during training. TLDR outperforms all prior works, except in HalfCheetah compared to METRA. METRA learns low-dimensional skills and extends the temporal distance along a few directions specified by the skills, providing a strong inductive bias for simple locomotion tasks like HalfCheetah. On the other hand, TLDR achieves much larger state coverage in complex environments than METRA, including AntMaze-Large, AntMaze-Ultra, and Quadruped-Escape, where all other methods struggle and only explore limited regions. This shows the strength of our method in the exploration of complex environments.
We then compare the goal-reaching performance of our method with PEG and METRA in Figure 5. We first report the average distance between goals and the last states of trajectories. The results in Figures 5(a), 5(b) and 5(c) show that TLDR can navigate towards the given goals closer than, or at least on par with METRA. For the AntMaze environments, we report the number of pre-defined goals reached by the goal-conditioned policy. Figures 5(d) and 5(e) show that TLDR is the only method that can navigate towards a various set of goals in both mazes, demonstrating its superior exploration and goal-conditioned policy learning with temporal distance.
@@
Figure 6: Results in pixel-based environments. We compare TLDR with prior works in pixel-based Quadruped and Kitchen environments. In Quadruped (Pixel), TLDR demonstrates a slow learning speed compared to METRA and LEXA. For Kitchen (Pixel), TLDR could interact with all six objects during training, but shows low success rates for evaluation.
Figure 6 shows the results in pixel-based environments. In Quadruped (Pixel), TLDR explores diverse regions but learns slower than LEXA and METRA. For Kitchen (Pixel), TLDR interacts with all six objects during training, but struggles at learning the goal-conditioned policy. We hypothesize that learning a temporal abstraction is more challenging with pixel observations, which may lead $\phi$ to encode erroneous temporal information. We leave more detailed analyses for future works.
@@
Figure 7: Goal-reaching ability in AntMaze-Ultra. TLDR can cover more goals compared to METRA and PEG.
4.3 Qualitative Results
We visualize the learned goal-reaching behaviors on the AntMaze-Ultra environment in Figure 7. TLDR can successfully reach both near and faraway goals in diverse regions. On the other hand, METRA and PEG fail to navigate to diverse goals. METRA could reach some goals distant from the initial position, whereas PEG fails to reach temporally faraway goals. This clearly shows the benefit of using temporal distance in unsupervised GCRL.
@@ 4.4 Ablation Studies
To investigate the importance of temporal distance-aware representations in our algorithm, we conduct ablation studies on exploration strategies and GCRL reward designs.
@@
For goal selection and exploration rewards, we replace temporal distance, $\lVert\phi(\mathbf{s})-\mathbf{z}^{(j)}\rVert$ in Equation 2, with other exploration bonuses: RND, APT (with ICM [37] representations), and Disagreement. Note that goal-conditioned policies are still trained with the same temporal distance-based rewards as TLDR, thereby comparing only exploration strategies. As shown in Figure 8(a), using TLDR reward for goal selection and exploration rewards achieves significantly higher performance than other exploration bonuses. This result implies that our temporal distance-based rewards are effective for unsupervised exploration.
GCRL reward design.
We compare with two goal-conditioned policy learning methods: (1) QRL [14], which uses a quasimetric value function, and (2) sparse HER [8], which uses the sparse goal-reaching reward $-\mathbbm{1}(\mathbf{s}\neq\mathbf{g})$. Figure 8(b) shows the superior performance of our temporal distance-based GCRL reward. This highlights the importance of incorporating temporal distance-aware representations in training goal-conditioned policies.
(a) TLDR with different exploration methods
(b) TLDR with different GCRL rewards
Figure 8: Impact of temporal distance-aware representations in exploration and GCRL reward design. We evaluate our method with different design choices for (a) exploration methods and (b) GCRL rewards on Ant and AntMaze-Large. TLDR shows better state coverages than its ablated versions in both ablation studies, indicating the importance of using temporal distance for both exploration and GCRL.
@@ 5 Conclusion
In this paper, we introduce TLDR, an unsupervised GCRL algorithm that incorporates temporal distance-aware representations. TLDR leverages temporal distance for exploration and learning the goal-reaching policy. By pursuing states with larger temporal distances, TLDR can continuously explore challenging regions, achieving better state coverage. The experimental results demonstrate that our method can cover significantly larger state spaces across diverse environments than existing unsupervised reinforcement learning algorithms.
5.1 Limitations
While TLDR achieves remarkable state coverage, it still has several limitations. Firstly, TLDR shows a slower learning speed compared to METRA in pixel-based environments. Secondly, our temporal distance-aware representations do not capture the asymmetric temporal distance between the states, which can make policy learning challenging for asymmetric environments. Finally, TLDR achieves high efficiency in terms of wall clock time, but with a relatively low update-to-data ratio (number of gradient steps divided by number of environment steps) of $1/32$ in state-based experiments, as used in METRA. However, we believe that increasing the update-to-data ratio or using model-based approaches could potentially enhance sample efficiency.
@@ Acknowledgments
This work was supported in part by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant (RS-2020-II201361, Artificial Intelligence Graduate School Program (Yonsei University)) and the National Research Foundation of Korea (NRF) grant (RS-2024-00333634), and the Electronics and Telecommunications Research Institute (ETRI) grant (24ZR1100) funded by the Korean Government (MSIT).
@@ A.1 Computing Resources and Experiments
All experiments are done on a single RTX 4090 GPU and $4$ CPU cores. Each state-based experiment takes $12$ hours for all methods, following METRA [1], which trains each method for $9$-$10$ hours. It corresponds to the different environment steps used for different experiments, as described in Table 1. We use $5$ random seeds for all experiments, and report the mean and standard deviation of the results.
@@ A.2 Implementation Details
where we apply affine-transformed softplus $f$ to Equation 1:
@@
which alleviates the effect of too long distances $\lVert\phi(\mathbf{s})-\phi(\mathbf{g})\rVert$, following QRL [14].
@@ A.3 Hyperparameters
For METRA, we use $2$-D continuous skills for Ant, $16$-D discrete skills for HalfCheetah, $24$-D discrete skills for Kitchen (Pixel), and $4$-D continuous skills for other environments. We use the batch size of $1024$ for state-based environments and $256$ for pixel-based environments. We set the number of gradient steps for each experiment to be the same as ours. We use the default values for the remaining hyperparameters. To perform goal-reaching tasks with METRA, we set the skill $\mathbf{z}$ as $\frac{\phi(\mathbf{g})-\phi(\mathbf{s})}{\lVert\phi(\mathbf{g})-\phi(\mathbf{s})\rVert}$ for continuous skills or $\arg\max_{\text{dim}}\left(\phi(\mathbf{g})-\phi(\mathbf{s})\right)$ for discrete skills.
In PEG, we use the same hyperparameters used in their AntMaze experiments. Since PEG uses the normalized goal space, we measure the range of the observations and normalize the goal states according to the minimum and maximum range.
@@
For APT with ICM encoder, RND, and Disagreement, we use the same hyperparemeters as in URLB [39].
For the ablation with QRL, we use the learning rate of $0.0003$ for critics. We use an (input dim)-$1024$-$1024$-$128$ network for the encoder, $256$-$1024$-$2048$ for the projector, IQE-maxmean head of $64$ components of size $32$, and $128$-$1024$-$128$ for the latent dynamics model. The transition loss is weighted by $1$. For HER, we use the discount factor $\gamma=0.99$.
@@
SAC entropy coefficient $0.01$ (Kitchen), target entropy as $(-\dim\mathcal{A})$ (others)
@@
# of gradient steps per epoch $50$ (Ant, HalfCheetah, Humanoid-Run, Quadruped-Escape), $75$ (AntMaze-Large), $100$ (Kitchen), $150$ (AntMaze-Ultra), $200$ (Quadruped (Pixel))
@@ A.4 Environment Details
Quadruped-Escape is included in DeepMind Control Suite [32]. The quadruped robot is initialized in a basin surrounded by complex terrains, as described in Figure 3(d). Due to the complex terrains, moving further away from the initial position is challenging. Similar to the AntMaze environments, we fix the terrain shape. Also, we add the global $x,y,z$ coordinates of the agent to the observation. Quadruped-Escape has $104$-D observation space with $12$-D action space. The episode length is $200$.
@@
We use antmaze-large-play-v2 in D4RL [33]. The observation and action spaces are the same as the Ant environment. The episode length is $300$. To make exploration more challenging, we fix the initial location of the agent to be the bottom right corner of the maze, as shown in Figure 3(e).
@@
We use antmaze-ultra-play-v0 proposed by Jiang et al. [34]. The observation and action spaces are the same as the Ant environment. The episode length is $600$. Similar to AntMaze-Large, we fix the initial location of the agent to be the bottom right corner of the maze, as shown in Figure 3(f).
@@
We use the pixel-based version of the Quadruped environment [32] used in METRA [1]. Specifically, we use the image size of $64\times 64\times 3$ with $200$ episode length.
@@
We use the pixel-based version of the Kitchen environment [42] used in METRA [1] and LEXA [10]. Specifically, we use the image size of $64\times 64\times 3$ with $50$ episode length. The action space has $9$ dimensions.

### v2

@@ 3.5 Learning Exploration Policy
Similar to LEXA [10], we alternate between goal-reaching episodes and exploration episodes. For goal-reaching episodes, we execute the goal-conditioned policy until the end of the episodes. For exploration episodes, we sample the timestep $T_{G}\sim\text{Unif}(0,T-1)$ at the beginning of each episode and execute the exploration policy if the current timestep $t\geq T_{G}$.
@@ 3.6 Learning Goal-Conditioned Policy
4 Experiments
In this paper, we propose TLDR, a novel unsupervised GCRL method that utilizes temporal distance-aware representations for both exploration and optimizing a goal-conditioned policy. Through our experiments, we aim to answer the following three questions: (1) Does TLDR explore better compared to other exploration methods? (2) Is our goal-conditioned policy better than prior unsupervised GCRL methods? (3) How crucial is TLDR for goal-conditioned policy learning and exploration?
4.1 Experimental Setup
Ant
HalfCheetah
Humanoid-Run
Quadruped-Escape
AntMaze-Large
AntMaze-Ultra
Figure 3: We evaluate our method on $6$ state-based robotic locomotion environments.
@@ 4.1 Experimental Setup
Figure 4: State coverage in state-based environments. We measure the state coverage of unsupervised exploration methods. Our method consistently shows superior state coverage compared to other methods, except in HalfCheetah compared against METRA.
@@
As illustrated in Figure 3, we evaluate TLDR in $6$ state-based environments: Ant and HalfCheetah from OpenAI Gym [30], Humanoid-Run and Quadruped-Escape from DeepMind Control Suite (DMC) [31], AntMaze-Large from D4RL [32], and AntMaze-Ultra [33]. For Humanoid-Run and Quadruped-Escape, we include the 3D coordinates of the agents in their observations. In addition, we also evaluate on two pixel-based environments: Quadruped (Pixel) from METRA [1] and Kitchen (Pixel) from D4RL [32], with the $64\times 64\times 3$ image observation.
@@
  • METRA [1]: leverages temporal distance-aware representations for skill discovery.
  • PEG [2]: plans to obtain goals with maximum exploration rewards.
@@
(a) Ant
(b) HalfCheetah
(c) Humanoid-Run
(d) AntMaze-Large
(e) AntMaze-Ultra
Figure 5: Goal-reaching metrics of a goal-conditioned policy. For (a) Ant, (b) HalfCheetah, and (c) Humanoid-Run, we report the average distance between goals and the last states of trajectories (lower is better). TLDR achieves a comparable average goal distance to METRA. For AntMaze environments, we report the number of pre-defined goals reached by a goal-reaching policy ($7$ for (d) AntMaze-Large and $21$ for (e) AntMaze-Ultra), and TLDR significantly outperforms prior works.
@@ 4.2 Quantitative Results
In Figure 4, we compare the state coverage during training (i.e. the number of $1\times 1$ sized $(x,y)$-bins occupied by any of the training trajectories). TLDR outperforms all prior works, except in HalfCheetah compared to METRA. METRA learns low-dimensional skills and focuses on extending the temporal distance along a few directions specified by the skills, providing a strong inductive bias for simple locomotion tasks like HalfCheetah. On the other hand, TLDR achieves much larger state coverage in complex environments than METRA, including AntMaze-Large, AntMaze-Ultra, and Quadruped-Escape, where all other methods struggle and only explore limited regions. This shows the strength of our method in the exploration of complex environments.
We then compare the goal-reaching performance of TLDR with PEG and METRA in Figure 5 by measuring the average distance between goals and the last states of trajectories. The results in Figures 5(a), 5(b) and 5(c) show that TLDR can navigate towards the given goals closer than or at least on par with METRA. Figures 5(d) and 5(e) show that TLDR is the only method that can navigate towards a various set of goals in both mazes, demonstrating its superior exploration and goal-conditioned policy learning with temporal distance.
In Appendix B, we further show the comparisons in environment steps, not in hours. PEG shows better sample efficiency in relatively low-dimensional or easy-exploration tasks, such as Ant and HalfCheetah. However, the state coverages of PEG quickly converge to narrower regions, especially in AntMazes, than those of TLDR. METRA generally shows worse sample efficiency than TLDR.
@@
Figure 6: Results in pixel-based environments. We compare TLDR with prior works in pixel-based Quadruped and Kitchen environments. In Quadruped (Pixel), TLDR demonstrates a slow learning speed compared to METRA and LEXA. For Kitchen (Pixel), TLDR could interact with all six objects during training but shows low success rates for evaluation.
Figure 6 shows the results in pixel-based environments. In Quadruped (Pixel), TLDR explores diverse regions but learns slower than LEXA and METRA. For Kitchen (Pixel), TLDR interacts with all six objects during training, but struggles at learning the goal-conditioned policy. Further analysis in Appendix F suggests that the performance bottleneck is related to goal-conditioned policy learning with pixel observations. We leave more detailed analyses for future works.
4.3 Qualitative Results
@@
Figure 7: TLDR can cover more goals compared to METRA and PEG in AntMaze-Ultra.
Figure 7 visualizes the learned goal-reaching behaviors on the AntMaze-Ultra environment. TLDR can successfully reach both near and faraway goals in diverse regions. On the other hand, METRA and PEG fail to navigate to diverse goals. METRA could reach some goals distant from the initial position, whereas PEG fails to reach temporally faraway goals. This clearly shows the benefit of using temporal distance in unsupervised GCRL. More qualitative results can be found in Appendix D.
@@ 4.4 Ablation Studies
To investigate the importance of temporal distance-aware representations in our algorithm, we conduct ablation studies on GCRL reward designs and exploration strategies.
GCRL reward design.
We compare with three different goal-conditioned policy learning methods: (1) QRL [14], which learns a quasimetric value function and latent dynamics model, (2) sparse HER [8], which uses the sparse goal-reaching reward $-\mathbbm{1}(\mathbf{s}\neq\mathbf{g})$, and (3) DDL [19], which uses expected temporal distances as rewards. Figure 8(a) and Figure 21 show the superior performance of our temporal distance-based GCRL reward over HER and DDL, suggesting the importance of using optimal temporal distance as a dense reward signal. Furthermore, although QRL learns a value function that preserves optimal temporal distances, it struggles to learn an effective goal-reaching policy. Unlike QRL, which directly uses the learned value function along with an additional latent dynamics model, TLDR leverages temporal distance-aware representations to compute dense rewards for the goal-conditioned policy and shows better performance. In Appendix G, we show that this trend also holds with a fixed dataset, which ignores the effect of exploration and only compares goal-reaching reward designs for goal-reaching performances.
(a) TLDR with different GCRL rewards
(b) TLDR with different exploration methods
Figure 8: We evaluate our method with different design choices for (a) GCRL rewards and (b) exploration methods on Ant and AntMaze-Large. TLDR shows better state coverages than its ablated versions in both ablation studies, indicating the importance of using temporal distance-aware representations for both exploration and GCRL.
@@
For goal selection and exploration rewards, we replace TLDR reward in Equation 2, with other exploration bonuses: APT (with ICM [36] representations), RND, and Disagreement. Note that goal-conditioned policies are still trained with the same temporal distance-based rewards as TLDR, thereby comparing only exploration strategies. As shown in Figure 8(b), using TLDR reward for goal selection and exploration rewards achieves significantly higher performance than other exploration bonuses. This result implies that our temporal distance-based rewards are effective for unsupervised exploration.
@@ 5 Conclusion
In this paper, we introduce TLDR, an unsupervised GCRL algorithm that incorporates temporal distance-aware representations. TLDR leverages temporal distance for exploration and learning the goal-reaching policy. By pursuing states with larger temporal distances, TLDR can continuously explore challenging regions, achieving better state coverage. The experimental results demonstrate that TLDR can cover significantly larger state spaces across diverse environments than existing unsupervised RL algorithms.
Limitations.
While TLDR achieves remarkable state coverages, it still has several limitations:
  • TLDR shows a slow learning speed compared to METRA in pixel-based environments. Our analysis in Appendix F demonstrates the need for further research in representation learning and GCRL for pixel observations.
  • Our temporal distance-aware representations do not capture asymmetric temporal distance between states, which can make policy learning challenging for highly asymmetric environments.
  • Applying unsupervised RL to real robots has many challenges, including safety. While not tested on real robots, our preliminary results in Appendix E indicate that combining TLDR with safety-aware techniques [37, 38] is a promising future direction for real robotic systems.
  • TLDR achieves high efficiency in terms of wall clock time, but not in terms of sample efficiency, as shown in Appendix B. We believe that increasing the update-to-data ratio or using model-based RL could enhance the sample efficiency of our method.
@@ Acknowledgments
This work was supported in part by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant (RS-2020-II201361, Artificial Intelligence Graduate School Program (Yonsei University)), the National Research Foundation of Korea (NRF) grant (RS-2024-00333634), and the Electronics and Telecommunications Research Institute (ETRI) grant (24ZR1100) funded by the Korean Government (MSIT).
@@ A.1 Computing Resources and Experiments
All experiments are done on a single RTX 4090 GPU and $4$ CPU cores. Each state-based experiment takes $12$ hours for all methods, following METRA [1], which trains each method for $10$-$12$ hours. We report the number of environment steps used for the methods in our experiments in Table 1. We use $5$ random seeds for all experiments and report the mean and standard deviation of the results.
@@ A.2 Implementation Details
where $f$ is an affine-transformed softplus function:
@@
which prevents the distances $\lVert\phi(\mathbf{s})-\phi(\mathbf{g})\rVert$ from diverging, following QRL [14].
@@ A.3 Hyperparameters
For METRA, we use $2$-D continuous skills for Ant, $16$-D discrete skills for HalfCheetah, $24$-D discrete skills for Kitchen (Pixel), and $4$-D continuous skills for other environments. We use the batch size of $1024$ for state-based environments and $256$ for pixel-based environments. We set the number of gradient steps per epoch for each experiment to be the same as ours. We use the default values for the remaining hyperparameters. To perform goal-reaching tasks with METRA, we set the skill $\mathbf{z}$ as $\frac{\phi(\mathbf{g})-\phi(\mathbf{s})}{\lVert\phi(\mathbf{g})-\phi(\mathbf{s})\rVert}$ for continuous skills or $\arg\max_{\text{dim}}\left(\phi(\mathbf{g})-\phi(\mathbf{s})\right)$ for discrete skills.
In PEG, we use the same hyperparameters used in their ntMaze experiments. Since PEG uses the normalized goal space, we measure the range of the observations and normalize the goal states according to the minimum and maximum range.
@@
For APT (with ICM encoder), RND, and Disagreement, we use the same hyperparameters as in URLB [40].
For the ablation with QRL, we use the learning rate of $0.0003$ for the critic. We use an (input dim)-$1024$-$1024$-$128$ network for the encoder, $256$-$1024$-$2048$ for the projector, IQE-maxmean head of $64$ components of size $32$, and $128$-$1024$-$1024$-$128$ for the latent dynamics model. The transition loss is weighted by $1$. For HER, we use the discount factor $\gamma=0.99$.
@@
SAC entropy coefficient $0.01$ (Kitchen), target entropy as $(-\dim\mathcal{A})/2$ (others)
@@
# of gradient steps per epoch $50$ (Ant, HalfCheetah, Humanoid-Run, Quadruped-Escape), $75$ (AntMaze-Large), $100$ (Kitchen (Pixel)), $150$ (AntMaze-Ultra), $200$ (Quadruped (Pixel))
@@ A.4 Environment Details
Quadruped-Escape is included in DeepMind Control Suite [31]. The quadruped robot is initialized in a basin surrounded by complex terrains. Due to the complex terrains, moving further away from the initial position is challenging. Similar to the AntMaze environments, we fix the terrain shape. Also, we add the global $x,y,z$ coordinates of the agent to the observation. Quadruped-Escape has $104$-D observation space with $12$-D action space. The episode length is $200$.
@@
We use antmaze-large-play-v2 in D4RL [32]. The observation and action spaces are the same as the Ant environment. The episode length is $300$. To make exploration more challenging, we fix the initial location of the agent to be the bottom right corner of the maze, as shown in Figure 3 (AntMaze-Large).
@@
We use antmaze-ultra-play-v0 proposed by Jiang et al. [33]. The observation and action spaces are the same as the Ant environment. The episode length is $600$. Similar to AntMaze-Large, we fix the initial location of the agent to be the bottom right corner of the maze, as shown in Figure 3 (AntMaze-Ultra).
@@
We use the pixel-based version of the Quadruped environment [31] used in METRA [1]. Specifically, we use the image size of $64\times 64\times 3$ with episode length of $200$.
@@
We use the pixel-based version of the Kitchen environment [43] used in METRA [1] and LEXA [10]. Specifically, we use the image size of $64\times 64\times 3$ with the episode length of $50$. The action space has $9$ dimensions.